OpenStudy (annie96):

Use the formula for the area of a rectangle A = lw to answer the question. A rectangle has a width of n units. Its length is 4 units longer than the width. What is the area of the rectangle? (Points : 1) 2n + 4 2 + 4n 4n n(n + 4)

OpenStudy (anonymous):

You can plug in random numbers to solve this if you want.

OpenStudy (anonymous):

Say that you choose n = 2 l = 2 + 4 A = lw = 6(2) = 12 So try to plug them into these answer choices.

OpenStudy (annie96):

okay thank you